Please help me solve this equation:
What is the standard form of the equation of the line that is perpendicular to 6x+5y=0 and contains (-6,-1)?

Answer by mananth(16946)

6 x + 5 y = 0
Find the slope of this line

5 y = -6 x + 0
Divide by 5
y = -1.200 x + 0
Compare this equation with y=mx+b
slope m = -1.20

The slope of a line perpendicular to the above line will be the negative reciprocal 0.83
m1*m2=-1
The slope of the required line will be 5/6

m= 5/6 ,point ( -6 , -1 )
Find b by plugging the values of m & the point in
y=mx+b
-1 = -5.00 + b
b= 4
m= 5/6
The required equation is y = 5/ 6 x + 4
m.ananth@hotmail.ca